OpenStudy (anonymous):

my last question that i cant miss , please help. 5 8/9 divided by 8 2/3 = ?

OpenStudy (anonymous):

plug it in the calculator, do you have 1?

OpenStudy (anonymous):

The easiest way (IMO) for this sort of problem is to turn everything into proper fractions. For example, in small steps: \[5 \frac{8}{9}\] is the same as \[5 + \frac{8}{9}\] which is the same as \[ \frac{5}{1} + \frac{8}{9}\] To add these, you'll need the same denominator in both fractions, so write as \[ \frac{5*9}{1*9} + \frac{8}{9}\] which is \[ \frac{45}{9} + \frac{8}{9}\] or \[ \frac{53}{9}\] Once each term is a proper fraction, dividing them should be a piece of cake :)